# Observation Interference in Partially Observable Assistance Games

### v1

@@ 8 Conclusion
Even when the AI assistant and the human have perfect value alignment, we show how observation interference can emerge from several distinct incentives. As we focus on optimal assistants—analyzing optimal policy pairs and best responses—all of the incentives for observation interference that we consider are done for the human’s benefit. This creates a nuanced picture, suggesting that not all observation interference is inherently bad. In practice, we expect that AI assistants will exhibit observation interference for a mix of good and bad reasons. With this theory, our goal is to lay a foundation for understanding the causes of observation interference and helping to disentangle them in practice.
@@
We choose to study optimal solutions, such as optimal policy pairs and best responses. This has the advantage of providing general insight into the underlying game structure that is independent of any particular learning algorithm. However, this independence is also a drawback; if algorithms fail to find optimal solutions, they might break down in unexpected ways not captured by our theory.
Moreover, we find that some optimal policy pairs in our examples, such as Example 5.1, require $\mathbf{H}$ and $\mathbf{A}$ to have a shared communication protocol. It would be interesting to study additional solution concepts, such as correlated equilibria and communication equilibria, to handle this sort of communication [13]. While we consider only a single human and single assistant, it would also be interesting to study scenarios with multiple humans and multiple assistants. Lastly, while we run experiments in one model of a POAG, it would be interesting to see if and how our experimental trends generalize to other POAGs.
@@ Acknowledgments
Our techniques are similar to those of Shah et al. [31] and Desai [11], who show how to form a single-agent POMDP for $\mathbf{A}$ by embedding $\mathbf{H}$ into the environment dynamics. However, our construction works in the opposite direction, with $\mathbf{H}$ embedding $\mathbf{A}$’s actions and observations into the environment.
@@ B.1 A lemma about policies with internal states
The result is related to Kuhn’s [20] proof of the equivalence of behavioral and mixed strategies in perfect-recall extensive-form games.
@@
  • $\mathbf{H}$’s observations are given as follows. If $s\in\{E,I\}$ or $s_{0}\neq 1$, then $\mathbf{H}$ simply observes null. If $s\notin\{E,I\}$ and $s_{0}=1$, then $O^{\mathbf{H}}(o^{\mathbf{H}}|s,a^{\mathbf{A}},a^{\mathbf{H}})=\mathbb{1}[o^{\mathbf{A}}_{i}{=}s_{i+1}a^{\mathbf{A}}_{i}]$. That is, for each availability bit, $\mathbf{H}$ observes 0 if $\mathbf{A}$ set the availability bit to 00; otherwise, $\mathbf{H}$ simply observes the availability bit.
  • $R(s,a^{\mathbf{H}},a^{\mathbf{A}})=0$ if $s\in\{E,I\}$ or $s_{0}=0$. Otherwise, $R(s,a^{\mathbf{H}},a^{\mathbf{A}})=s_{a^{\mathbf{H}}}s_{a^{\mathbf{H}}+10}$. That is, a reward of $1$ is obtained if and only if the cuda version chosen by $\mathbf{H}$ is both available and compatible.
@@
Next, note that the only non-interference action for $\mathbf{A}$ is $(1,1,...,1)$. Thus, the only non-interference policy for $\mathbf{A}$ is to always play $(1,1,...,1)$. Call this policy $\pi^{\mathbf{A}}_{\text{ni}}$.
@@
Roughly, take any deterministic optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$. Consider the assistant policy ¯$\bar{}\pi^{\mathbf{A}}$ that at each time step communicates $\mathbf{A}$’s full observation to $\mathbf{H}$ and that replaces interference with non-interference actions. Because $\pi^{\mathbf{A}}$ is deterministic, $\mathbf{H}$ can infer what $\pi^{\mathbf{A}}$ would have communicated based on ¯$\bar{}\pi^{\mathbf{A}}$’s communications. The rest of the proof goes the same way as Theorem 4.2. ∎
@@
For the proof of Theorem 4.7, we’ll use the concept of entropy. For any probability distribution $P$ over some discrete space, let $H(P)\coloneqq-\sum_{x}P(x)\log P(x)$ denote the distribution’s entropy. The following is a well-known result in information theory [e.g., 2, Theorem 1.4.5; 9, Theorem 2.6.5].
@@
Let $L$ be some set of states. Let $(P_{a}(\cdot\mid s))_{s\in L}$ and $(P_{b}(\cdot\mid s))_{s\in L}$ be families of probability distributions s.t. $P_{a}$ is strictly more informative than $P_{b}$ with transformation function $F$. Further let $S$ be some random variable over $L$ with full support. Let $X_{a}\sim P_{a}(\cdot\mid S)$ and $X_{b}\sim F(X_{a})$. Then $S$ and $X_{a}$ are dependent given $X_{b}$. In particular, from Lemma B.3 we get that $\mathbb{E}_{X}\left[H(P(S\mid X))\right]<\mathbb{E}_{\hat{X}}\left[H(P(S\mid\hat{X}))\right]$.
@@
Note that the entropy-minimizing policy used in the proof may still interfere with observations at the action level. For example, by default $\mathbf{H}$ might receive a low-information signal about the world. The entropy-minimizing policy might be one in which $\mathbf{A}$ overwrites this default signal in a way that expresses more information about the world. For instance, let’s assume that by default, $\mathbf{H}$ observes a random number between $-20$ and 00 if it’s cold outside and a random number between 00 and $+40$ if it’s warm outside. $\mathbf{A}$ receives various hints about the temperature and can overwrite the signal with an arbitrary number. (I.e., for each number between $-20$ and $+40$, there’s an action that sets $\mathbf{H}$’s observation to be that number.) Assuming nothing else happens in this POAG, the entropy-minimizing policies will be ones that overwrite the signal in a way that encodes $\mathbf{A}$’s information about the temperature. For instance, $\mathbf{A}$ it may (or may not) be an non-interfering-at-the-policy-level strategy for $\mathbf{A}$ to overwrite $\mathbf{H}$’s signal with $\mathbf{A}$’s expectation of the temperature in degrees Celsius. Given such a policy, the entropy of $\mathbf{H}$’s beliefs about the world is lower than before ($\mathbf{H}$ has more information about the temperature). But each of these overwriting actions individually is observation-interfering.
@@
Suppose the utilities of the human choice are given as follows: $1$ for the favored CPU-optimized type; $1$ for a GPU-optimized node if $\mathbf{H}$ favors the GPU-optimized node. The reward is 00 otherwise. On the second step, the reward for the favored type of node is $10$ and 00 for the other type of node.
@@
    $\mathbb{1}[a^{\mathbf{H}}=1]\mathbb{1}[s\in\{s_{a},s_{b}\}]+\mathbb{1}[a^{\mathbf{H}}=2]\mathbb{1}[s\in\{s_{c},s_{d}\}]$. All other rewards are 00.
@@
Now consider the non-interference policy for $\mathbf{A}$ that always plays man. Then a Boltzmann-rational $\mathbf{H}$ will choose as follows: If she observes $s_{a}$ or $s_{c}$, then she will choose an expected utility of $7$ with probability $\propto\exp(7)$ and an expected utility of 00 with probability $\propto\exp(0)$. Thus, the expected utility is
@@
Now consider the interference policy for $\mathbf{A}$ in which $\mathbf{A}$ always plays tldr. Then upon observing either 00 or $1$, the human chooses between a utility of 00 and a utility of $4$. Thus, the expected utility is
$4\cdot\frac{\exp(4)}{\exp(4)+\exp(0)}\approx 3.92806$ (8)
@@
Let $k-1$ be the $k$-th best action according to 00 and let $k$ be the $k+1$-th best action according to 00. By the tower rule (Eq. 9), $y_{0,k-1}-y_{0,k}$ must be greater than $y_{s,k-1}-y_{s,k}$ for some $s$. (If the difference in these expected utilities changes when the signal is observed, then it must sometimes decrease.) But then in cases where this difference decreases as $s$ is observed, we clearly have that the difference between one of the $k$ best actions to the $k+1$-th best action under $s$ also decreases. ∎
@@
WLOG let 00 be the best action under all signals, $1$ the second-best and so on. Let $k$ be the largest number that the differences between the utilities of actions $10,...,k-$ 1 are always the same. (Typically $k=0$.) Let $\tilde{S}$ be the set of signals under which the difference to the utility of $k$ (the $k+1$-th best action) is minimized. Note that in particular, the difference must be smaller than under 00 by Lemma E.3. WLOG assume that for all signals, $k$ is among the $k+1$-th best actions.
@@
We first provide a bound on $A$. We first show that $A<0$. To show this, note first that in all enumerators in $A$, we can replace $y_{\tilde{s},a}$ with $y_{0,a}$ (by choice of $\tilde{s}$ and $k$). So all we need to show is that the second denominator is smaller than the first, i.e., $\sum_{a^{\prime}}\exp(\beta y_{\tilde{s},a^{\prime}})>\sum_{a^{\prime}}\exp(\beta y_{0,a^{\prime}})$. But this this is easy to see from the fact that $y_{\tilde{s},a}=y_{0,a}$ for $a=0,1...,k-1$ and $y_{\tilde{s},k}>y_{0,k}$. For large $\beta$, $\exp(\beta y_{\tilde{s},k})$ will be much larger than $\sum_{a^{\prime}=k,k+1,...}\exp(\beta y_{0,a^{\prime}})$.
@@
Next we upper bound $B$. First, the best case for the effect on … is that all the probability mass that under 00 is on the top $k$ actions ends up on the $k$-th best action, i.e.,
@@
Finally, we must upper bound $C$. First, we can upper bound $C$ by considering a case where all probability mass that in 00 was outside the top $k$ actions, goes to the best action when a signal outside of $\tilde{S}$ is observed, i.e.,
@@ Appendix F Proof of $\mathbf{A}$’s best response in the product selection game
Consider $\mathbf{A}$’s perspective. $\mathbf{A}$’s interference is equivalent to selecting a set of $d-k$ untampered products from which $\mathbf{H}$ selects according to a Boltzmann distribution on $H_{i}$. As $\mathbf{A}$ neither sees nor affects the $H_{i}$, by symmetry, over all draws of the game, $\mathbf{H}$ selects each of the $d-k$ products with equal probability. $\mathbf{A}$’s expected payoff for choosing $d-k$ products, then, is the uniform average of the products’ expected $U_{i}$.
How does $\mathbf{A}$ choose the set of $d-k$ products to maximize the uniform average of the products’ expected $U_{i}$? Recall $U_{i}=H_{i}+R_{i}$. As $\mathbf{A}$ neither sees nor affects the $H_{i}$, $\mathbf{A}$ can ignore the $H_{i}$ and consider only the $R_{i}$. Denote the expected $R_{i}$ by $\hat{R}_{i}=\mathbb{E}[R_{i}]$. If $\mathbf{A}$ observes $R_{i}$, then $\hat{R}_{i}=R_{i}$. If $\mathbf{A}$ doesn’t observe $R_{i}$, then $\hat{R}_{i}=0.5$. To choose the maximum $d-k$ values for $\hat{R_{i}}$, $\mathbf{A}$ interferes with the minimum $k$ values of $\hat{R}_{i}$. ∎
@@ Appendix H Code assets
Our experiments use the Python software libraries Matplotlib [18], NumPy [16], pandas [27, 36], and seaborn [34].

### v2

@@ 8 Conclusion
Optimal policy pairs sometimes require $\mathbf{H}$ and $\mathbf{A}$ to have a shared communication protocol (e.g., Example 5.1). It would be interesting to study additional solution concepts, such as correlated equilibria and communication equilibria, to handle this sort of communication (Forges, 1986). While we consider only a single human and single assistant, it would also be interesting to study scenarios with multiple humans and multiple assistants. As we focus on the Boltzmann rationality model of human decision making (Definition 6.1), future work could consider other human models and empirical validation with human subjects. Lastly, while we run experiments in one model of a POAG, it would be interesting to see if and how our experimental trends generalize to other POAGs, including POAGs where $\mathbf{A}$ must query $\mathbf{H}$’s preferences.
Impact Statement
AI assistants are being developed and deployed in settings where humans can only partially observe what’s happening. For example, AI assistants including ChatGPT, Claude, and Gemini can search the web while only returning summaries to users (OpenAI, 2024; Anthropic, 2025; Google, 2024). Moreover, the sorts of AI models powering these assistants are processing increasingly long inputs. Whereas the original ChatGPT model could only process 4096 input tokens, today’s Gemini 1.5 Pro can process 2,000,000 input tokens—which is roughly 100,000 lines of code, or 16 novels of average length in English (Google, 2025). In the future, we anticipate that AI assistants will be deployed at increasing scale, independently taking more actions on behalf of users and processing increasingly long context lengths. We thus expect that over time, humans will have less and less ability to directly observe everything that’s happening.
Even when the AI assistant and the human have perfect value alignment, we show how observation interference can emerge from several distinct incentives. As we focus on optimal assistants—analyzing optimal policy pairs and best responses—all of the incentives for observation interference that we consider are done for the human’s benefit. This creates a nuanced picture, suggesting that not all observation interference is inherently bad. As AI assistants might exhibit observation interference for a mix of good and bad reasons, it would be interesting for future work to explore how to handle this nuanced situation. For example, future AI systems could be designed with transparency about when interference occurs and user controls to override interference when desired.
With this theory, our goal is to understand the causes of observation interference and help disentangle them in practice. We intend for our work to help AI developers build assistants that their users can trust. Our work is primarily theoretical, and we are not aware of any ways it could be used to cause harm.
Computational Complexity
Given that finding optimal policies in POAGs is NEXP-hard, how might our results apply in a given environment? Most of our paper is descriptive, characterizing when observation tampering could happen. Complexity considerations could affect these results in either direction. It’s easy to construct environments where finding good observation-interfering policies is computationally intractable but constructing good non-interfering policies is easy; and vice versa. In practice, complexities of the environment can be orthogonal to incentives to interfere. For instance, a real-world version of the CUDA example is complex (A assesses complicated software compatibility issues), but the decision whether to interfere with observations is easy. We believe our characterizations remain useful even in complex environments (where we can’t expect optimal policies), although we can’t make as definitive claims as we can about optimal policies.
We have discussed allowing communication between H and A. A complexity-theoretic argument favors this solution: If H and A share all private information, the game effectively turns from a DecPOMDP into a POMDP. Solving POMDPs is PSPACE-complete and thus likely easier than solving DecPOMDPs.
@@ Acknowledgments
Our techniques are similar to those of Shah et al. (2020) and Desai (2017), who show how to form a single-agent POMDP for $\mathbf{A}$ by embedding $\mathbf{H}$ into the environment dynamics. However, our construction works in the opposite direction, with $\mathbf{H}$ embedding $\mathbf{A}$’s actions and observations into the environment.
@@ B.1 A Lemma about Policies with Internal States
The result is related to Kuhn’s (Kuhn, 1953) proof of the equivalence of behavioral and mixed strategies in perfect-recall extensive-form games.
@@
  • $\mathbf{H}$’s observations are given as follows. If $s\in\{E,I\}$ or $s_{0}\neq 1$, then $\mathbf{H}$ simply observes null. If $s\notin\{E,I\}$ and $s_{0}=1$, then $O^{\mathbf{H}}(o^{\mathbf{H}}|s,a^{\mathbf{A}},a^{\mathbf{H}})=\mathbb{1}[o^{\mathbf{A}}_{i}{=}s_{i+1}a^{\mathbf{A}}_{i}]$. That is, for each availability bit, $\mathbf{H}$ observes 0 if $\mathbf{A}$ set the availability bit to 0; otherwise, $\mathbf{H}$ simply observes the availability bit.
  • $R(s,a^{\mathbf{H}},a^{\mathbf{A}})=0$ if $s\in\{E,I\}$ or $s_{0}=0$. Otherwise, $R(s,a^{\mathbf{H}},a^{\mathbf{A}})=s_{a^{\mathbf{H}}}s_{a^{\mathbf{H}}+10}$. That is, a reward of 11 is obtained if and only if the cuda version chosen by $\mathbf{H}$ is both available and compatible.
@@
Next, note that the only non-interference action for $\mathbf{A}$ is (1,1,,1)(1,1,...,1). Thus, the only non-interference policy for $\mathbf{A}$ is to always play (1,1,,1)(1,1,...,1). Call this policy $\pi^{\mathbf{A}}_{\text{ni}}$.
@@
Roughly, take any deterministic optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$. Consider the assistant policy $\bar{}\pi^{\mathbf{A}}$ that at each time step communicates $\mathbf{A}$’s full observation to $\mathbf{H}$ and that replaces interference with non-interference actions. Because $\pi^{\mathbf{A}}$ is deterministic, $\mathbf{H}$ can infer what $\pi^{\mathbf{A}}$ would have communicated based on $\bar{}\pi^{\mathbf{A}}$’s communications. The rest of the proof goes the same way as Theorem 4.2. ∎
@@
For the proof of Theorem 4.7, we’ll use the concept of entropy. For any probability distribution $P$ over some discrete space, let $H(P)\coloneqq-\sum_{x}P(x)\log P(x)$ denote the distribution’s entropy. The following is a well-known result in information theory [e.g., 3, Theorem 1.4.5; 10, Theorem 2.6.5].
@@
[…]
@@
Note that the entropy-minimizing policy used in the proof may still interfere with observations at the action level. For example, by default $\mathbf{H}$ might receive a low-information signal about the world. The entropy-minimizing policy might be one in which $\mathbf{A}$ overwrites this default signal in a way that expresses more information about the world. For instance, let’s assume that by default, $\mathbf{H}$ observes a random number between 20-20 and 0 if it’s cold outside and a random number between 0 and +40+40 if it’s warm outside. $\mathbf{A}$ receives various hints about the temperature and can overwrite the signal with an arbitrary number. (I.e., for each number between 20-20 and +40+40, there’s an action that sets $\mathbf{H}$’s observation to be that number.) Assuming nothing else happens in this POAG, the entropy-minimizing policies will be ones that overwrite the signal in a way that encodes $\mathbf{A}$’s information about the temperature. For instance, $\mathbf{A}$ it may (or may not) be an non-interfering-at-the-policy-level strategy for $\mathbf{A}$ to overwrite $\mathbf{H}$’s signal with $\mathbf{A}$’s expectation of the temperature in degrees Celsius. Given such a policy, the entropy of $\mathbf{H}$’s beliefs about the world is lower than before ($\mathbf{H}$ has more information about the temperature). But each of these overwriting actions individually is observation-interfering.
@@
Suppose the utilities of the human choice are given as follows: $1$ for the favored CPU-optimized type; $1$ for a GPU-optimized node if $\mathbf{H}$ favors the GPU-optimized node. The reward is 0 otherwise. On the second step, the reward for the favored type of node is $10$ and 0 for the other type of node.
@@
    $\mathbb{1}[a^{\mathbf{H}}=1]\mathbb{1}[s\in\{s_{a},s_{b}\}]+\mathbb{1}[a^{\mathbf{H}}=2]\mathbb{1}[s\in\{s_{c},s_{d}\}]$. All other rewards are 0.
@@
Now consider the non-interference policy for $\mathbf{A}$ that always plays man. Then a Boltzmann-rational $\mathbf{H}$ will choose as follows: If she observes $s_{a}$ or $s_{c}$, then she will choose an expected utility of $7$ with probability $\propto\exp(7)$ and an expected utility of 0 with probability $\propto\exp(0)$. Thus, the expected utility is
@@
Now consider the interference policy for $\mathbf{A}$ in which $\mathbf{A}$ always plays tldr. Then upon observing either 0 or $1$, the human chooses between a utility of 0 and a utility of $4$. Thus, the expected utility is
$4\cdot\frac{\exp(4)}{\exp(4)+\exp(0)}\approx 3.92806.$ (8)
@@
Let k1k-1 be the $k$-th best action according to 0 and let $k$ be the $k+1$-th best action according to 0. By the tower rule (Eq. 9), $y_{0,k-1}-y_{0,k}$ must be greater than $y_{s,k-1}-y_{s,k}$ for some $s$. (If the difference in these expected utilities changes when the signal is observed, then it must sometimes decrease.) But then in cases where this difference decreases as $s$ is observed, we clearly have that the difference between one of the $k$ best actions to the $k+1$-th best action under $s$ also decreases. ∎
@@
WLOG let 0 be the best action under all signals, $1$ the second-best and so on. Let $k$ be the largest number that the differences between the utilities of actions 0,,k10,...,k-1 are always the same. (Typically $k=0$.) Let $\tilde{S}$ be the set of signals under which the difference to the utility of $k$ (the $k+1$-th best action) is minimized. Note that in particular, the difference must be smaller than under 0 by Lemma E.3. WLOG assume that for all signals, $k$ is among the $k+1$-th best actions.
@@
We first provide a bound on $A$. We first show that $A<0$. To show this, note first that in all enumerators in $A$, we can replace $y_{\tilde{s},a}$ with $y_{0,a}$ (by choice of $\tilde{s}$ and $k$). So all we need to show is that the second denominator is smaller than the first, i.e., $\sum_{a^{\prime}}\exp(\beta y_{\tilde{s},a^{\prime}})>\sum_{a^{\prime}}\exp(\beta y_{0,a^{\prime}})$. But this this is easy to see from the fact that $y_{\tilde{s},a}=y_{0,a}$ for a=0,1,k1a=0,1...,k-1 and $y_{\tilde{s},k}>y_{0,k}$. For large $\beta$, $\exp(\beta y_{\tilde{s},k})$ will be much larger than $\sum_{a^{\prime}=k,k+1,...}\exp(\beta y_{0,a^{\prime}})$.
@@
Next we upper bound $B$. First, the best case for the effect on … is that all the probability mass that under 0 is on the top $k$ actions ends up on the $k$-th best action, i.e.,
@@
Finally, we must upper bound $C$. First, we can upper bound $C$ by considering a case where all probability mass that in 0 was outside the top $k$ actions, goes to the best action when a signal outside of $\tilde{S}$ is observed, i.e.,
@@ Appendix F Proof of $\mathbf{A}$’s Best Response in the Product Selection Game
Consider $\mathbf{A}$’s perspective. $\mathbf{A}$’s interference is equivalent to selecting a set of dkd-k untampered products from which $\mathbf{H}$ selects according to a Boltzmann distribution on $H_{i}$. As $\mathbf{A}$ neither sees nor affects the $H_{i}$, by symmetry, over all draws of the game, $\mathbf{H}$ selects each of the dkd-k products with equal probability. $\mathbf{A}$’s expected payoff for choosing dkd-k products, then, is the uniform average of the products’ expected $U_{i}$.
How does $\mathbf{A}$ choose the set of dkd-k products to maximize the uniform average of the products’ expected $U_{i}$? Recall $U_{i}=H_{i}+R_{i}$. As $\mathbf{A}$ neither sees nor affects the $H_{i}$, $\mathbf{A}$ can ignore the $H_{i}$ and consider only the $R_{i}$. Denote the expected $R_{i}$ by $\hat{R}_{i}=\mathbb{E}[R_{i}]$. If $\mathbf{A}$ observes $R_{i}$, then $\hat{R}_{i}=R_{i}$. If $\mathbf{A}$ doesn’t observe $R_{i}$, then $\hat{R}_{i}=0.5$. To choose the maximum dkd-k values for $\hat{R_{i}}$, $\mathbf{A}$ interferes with the minimum $k$ values of $\hat{R}_{i}$. ∎
@@ Appendix H Code Assets
Our experiments use the Python software libraries Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), pandas (pandas development team, 2020; Wes McKinney, 2010), and seaborn (Waskom, 2021).